OpenStudy (anonymous):

what is the limit of (5+(-1)^n)/5 as n goes to infinity

OpenStudy (zarkon):

there is no 'x' in your expression so this should be easy ;)

OpenStudy (anonymous):

sorry, as n goes to zero

OpenStudy (anonymous):

the odd thing here is that although it is unnecessary, usually letters in the middle of the alphabet like \(i, j, k, l, m, n\) are re served for integers in any case there is some problem because if \(n\) can be any number, it could for example be \(\frac{1}{2}\) and then you would gave \((-1)^{\frac{1}{2}}=\sqrt{-1}\) so i think maybe there is some mistake

OpenStudy (anonymous):

perhaps it is the limit as \(n\to \infty\)

OpenStudy (anonymous):

oops, i'm sorry. It is as n goes to infinity..

OpenStudy (zarkon):

n going to zero is a much more interesting problem :)

OpenStudy (zarkon):

lrt letting n=1 then 2 then 3 then 4 .... see if you notice a pattern

OpenStudy (zarkon):

*try letting